OpenStudy (anonymous):

The radio station has placed pennies in a cylindrical glass jar. Each penny is 0.75 inches in diameter and 0.061 inches thick. If the cylindrical glass jar containing the pennies has a diameter of 6 inches and a height of 11.5 inches, how many pennies can fit inside the jar?

OpenStudy (anonymous):

@SolomonZelman I did the work, can you just see if it is correct? @shamim

OpenStudy (anonymous):

@Loser66 or you can check too :)

OpenStudy (loser66):

hahaha.. ok, show your work first

OpenStudy (anonymous):

Lol thanks! Hold on :)

OpenStudy (anonymous):

This is my answer because I have to show work and write a written response! Each Penny is 0.75 in diameter, and 0.065 inches thick. Cylindrical Glass jar is 6 inches in diameter, and 11.5 inches in height. Volume of cylindrical glass V = 3.14r^2 h V = 3.14 (3)^2 (11.5) V = 3.14 (9) (11.5) V = 3.14 (103.5) V = 324.99 inches. Volume of Pennies V = 3.14r ^2 h V = 3.14 (0.375)^2 (0.061) V = 0.4415625 ( 0.061) V = 0.026935312 Now divide the volume of the glass by the volume of the pennies. And round it to the nearest hundredths. 324.99 / 0.026935312 = 12065.57399 rounded to the nearest hundredths 12065.57

OpenStudy (anonymous):

Total of 985 pennies in the jar?

OpenStudy (loser66):

the top part is perfectly ok, but why total of 985 pennies? why is it not 12065 pennies?

OpenStudy (anonymous):

oooh I got the "totsl of 985 pennies" online. I guess it's wrong then. SO then 12,065 pennies would be my answer?

OpenStudy (anonymous):

total*

OpenStudy (loser66):

I think so. :)

OpenStudy (anonymous):

Yay, thank you!

OpenStudy (anonymous):

@SolomonZelman do you also agree? @Loser66 sorry I like to as two people lol

OpenStudy (anonymous):

ask

OpenStudy (loser66):

hahahah... It's Ok, friend.